OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation 8x – y = –9 4x – 3y = –22

OpenStudy (anonymous):

if u r solving for x and y then... 1st rearrange: 8x-y=9 becomes y=8x+9 (this is equation 1) and 4x-3y=-22 becomes y=(4x/3) + (22/3) (this is equation 2) rearrange equation 1 to get x=(y-9)/8 (this is equation 3) substitute equation 3 into 2: y= [4{(y-9)/8}/3] +(22/3) simplify: y= [(y-9)/6] +(22/3) bring [(y-9)/6] to the left of the equal sign y-[(y-9)/6]=(22/3) with calculator this'll equal 7 for working... (5y/6)+(3/2)=(22/3) start making y subject only...so (5y/6)=(22/3)-(3/2) (5y/6)=(35/6) 5y=(35/6)(6) 5y=25 y=7 then u substitute this value into first equation and solve for x......:) srry i made soo many steps

OpenStudy (anonymous):

;)...if i'm wrong someone correct me but it's unlikely...i checked with calculator